UMUC bank has a local subsidiary, which requires $75,000 cash every month. It uses $safe armored trucks service for cash delivery. Service company charges $500 for each delivery.
In order to send cash to its subsidiary, the bank has to liquidate part of its securities portfolio which generates 15% annual return.
How many armored truck service trips PER YEAR should the bank order?
(Hint: The annual rate of return in this question is given as 12 * actual monthly rate of return. You can round off to the nearest whole number,
for example if your answer is 17.62, you can answer 18.)
Answer #1
Maximum Optimum trip that should be undertaken in a mth.,
Equating the total delivery charges & the opportunity cost of loss of interest on $ 75000,
500*x=75000*15%/12
we get the no.of trips in a mth= 1.875
ie.1.875*12=
22.5
ie
23
Armored truck service trips PER YEAR the bank should order 23
Verification :
Delivery chgs. 22.5*500= 11250
loss of int. 22.5/1.875*75000*15%/12= 11250